OpenStudy (marcelie):

help please !!!! explain using theorems 4 5 7 and 9 , why the fucntion is continous at every number in its domain , state the domain. a(t) =arcsin(1+2t)

OpenStudy (marcelie):

@Directrix

OpenStudy (marcelie):

@mathmale

OpenStudy (marcelie):

@dan815

OpenStudy (anonymous):

who knows what the theorems are? every function you know is continuous on its domain the domain of arcsine is \[[-\frac{\pi}{2},\frac{\pi}{2}]\] set \[1+2t=-\frac{\pi}{2}\] and solve for \(t\) then do the same for \[\frac{\pi}{2}\] that will give you your domain

OpenStudy (marcelie):

well not really im so lost in this lesson

OpenStudy (marcelie):

why is it set to --pi/2?

OpenStudy (marcelie):

i dont understand :(

OpenStudy (anonymous):

because the domain of arcsine starts at \(-\frac{\pi}{2}\)

OpenStudy (marcelie):

oh but i got -3pi/4

OpenStudy (anonymous):

seems unlikely that if you solve \[1+2t=-\frac{\pi}{2}\] you get \[t=-\frac{3\pi}{4}\]

OpenStudy (marcelie):

yes i got tht

OpenStudy (anonymous):

you got what? \(-\frac{3\pi}{4}\)?

OpenStudy (marcelie):

yes.. idk what i did wrong

OpenStudy (anonymous):

takes two steps, just like solving \[1+2t=A\] for \(t\) subtract 1, divide by 2

OpenStudy (marcelie):

1/2

OpenStudy (anonymous):

\[1+2t=A\\ 2t=A-1\\t=\frac{A-1}{2}\]

OpenStudy (anonymous):

replace \(A\) by \(-\frac{\pi}{2}\)

OpenStudy (marcelie):



OpenStudy (marcelie):

@satellite73 come back

OpenStudy (marcelie):

@Directrix

Directrix (directrix):

What do these theorems state? The theorem numbers are different in different texts. >>theorems 4 5 7 and 9

OpenStudy (marcelie):

well im not sure im so confuse